# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 4.2. Quasi-convex approximation via supconvolution
Let $X\subset{\mathbb{R}}^{n}$ and let $u\in\operatorname{USC}(X)$ be bounded above. For each $\varepsilon>0$ the $\varepsilon$-supconvolution of $u$ the function defined by
@@
Before proving the results we need, here are three noteworthy facts.
@@
One can visualize how a function is transformed by the supconvolution operation by noticing that $u^{\varepsilon}$ is the upper envelope of all quadratic functions of radius $-2\varepsilon$ with vertex at some point of the graph of $u$ [^5] By upper envelope of a family of functions $\mathscr{F}$ we mean the function $g\vcentcolon=\sup_{f\in\mathscr{F}}f$. Also note that a quadratic function with negative radius is well-defined: it just opens downwards. Also, notice that, even if we decided to “preserve” the domain of $u$ in Definition 4.5 by defining its supconvolution only on $X$, $u^{\varepsilon}$ is in fact well-defined by (4.7) on the whole space ${\mathbb{R}}^{n}$.
@@
In Theorem 4.9 below, we prove that supconvolutions are quasi-convex functions. In the obvious manner, one can also define the inf-convolution and prove that it is quasi-concave. Moreover, it is not difficult to see that if $\varepsilon$ is sufficiently small, then the $\varepsilon$-supconvolution (resp. $\varepsilon$-inf-convolution) of a quasi-concave (resp. quasi-convex) function remains quasi-concave (resp. quasi-convex). This fact, along with the characterization given in Proposition 2.5, proves a part of a well-known result of Lasry and Lions [37] on sup-inf-convolutions being $C^{1,1}$.
Four important properties of the supconvolution are gathered in the following theorem. We have already partially revealed the first one.
@@
Let $u\in\operatorname{USC}(X)$ be bounded above, let $u^{\varepsilon}$ be its $\varepsilon$-supconvolution. The following hold:
@@
    $u^{\varepsilon}$ is $\frac{1}{\varepsilon}$-quasi-convex on ${\mathbb{R}}^{n}$;
@@
    on $X_{\delta}\vcentcolon=\{x\in X:\ d(x,\partial X)>\delta\}$.
@@
(i). It suffices to note that
$u^{\varepsilon}(\cdot)+\frac{1}{2\varepsilon}|\cdot|^{2}=\sup_{y\in{\mathbb{R}}^{n}}\bigg{(}u(y)-\frac{1}{2\varepsilon}|y|^{2}+\frac{1}{\varepsilon}\langle y,\cdot\rangle\bigg{)}$
@@
(ii). For each $x\in X$ fixed, one has
$u^{\varepsilon}(x):=\sup_{y\in X}v_{x}(y)\quad\text{where}\quad v_{x}(y):=u(y)-\frac{1}{2\varepsilon}|y-x|^{2}.$ (4.9)
Each function $v_{x}$ is upper semicontinuous on $X$ since it is the sum of upper semicontinuous functions. If $X\subsetneq{\mathbb{R}}^{n}$ and $X$ is closed, then extend each $v_{x}$ to be the upper semicontinuous function $\hat{v}_{x}$ defined on all of ${\mathbb{R}}^{n}$ by assigning the value $\infty$- ∞ on $X^{c}$ (see Remark 4.2(a)). Notice that with this extension, one can rewrite (4.9) as
$u^{\varepsilon}(x):=\sup_{y\in{\mathbb{R}}^{n}}\hat{v}_{x}(y),\quad\forall\,x\in{\mathbb{R}}^{n}.$ (4.10)
Moreover, since $\hat{v}_{x}\equiv\infty$ on $X^{c}$ and is defined by (4.9) on $X$ (where $u$ is bounded above), one has that $\hat{v}_{x}$ is anti-coercive . Hence, for each $x\in{\mathbb{R}}^{n}$ and $\varepsilon>0$ fixed, applying Proposition 4.3(d) to $\hat{v}_{x}$ gives the existence of a point $\xi=\xi(x,\varepsilon)\in{\mathbb{R}}^{n}$ which realizes $u^{\varepsilon}(x)$ given by (4.10).
(iii). Since $y=x\in X$ competes in the definition (4.7) of $u^{\varepsilon}(x)$ and since the penalization term in (4.7) is non positive, one has
$\forall\,x\in X:\quad u(x)\leqslant u^{\varepsilon^{\prime}}(x)\leqslant u^{\varepsilon}(x)\quad\text{whenever $\varepsilon^{\prime}\leqslant\varepsilon$}.$ (4.11)
@@
where we can write “$\max$” in place of “$\sup$” since both $u$ and its quadratic penalizations are upper semicontinuous and the closed balls are compact. It is easy to see that for each fixed $x$, $\rho(x,\varepsilon)\searrow 0$ as $\varepsilon\searrow 0$. Indeed, if not, then for all $\varepsilon$ sufficiently small, $\arg\max_{y\in X}\left(u-\frac{1}{2\varepsilon}|\cdot-\,x|^{2}\right)\subset B_{r}(x)^{\mathsf{c}}$, for some $r=r(x)>0$ independent of $\varepsilon$, yielding $u^{\varepsilon}(x)\leqslant\sup_{B_{r}(x)}u-\frac{r^{2}}{2\varepsilon}<u(x)$ if $\varepsilon$ is sufficiently small, which contradicts (4.11). Hence, the rightmost term in (4.12) converges to $u(x)$ as $\varepsilon\searrow 0$, [^6] The term converges to $\max\{u(x),\limsup_{y\to x}u(y)\}$, where $\limsup_{y\to x}u(y)\leqslant u(x)$ by the upper semicontinuity of $u$. yielding $u^{\varepsilon}(x)\searrow$ $u(x)$.
@@
(iv). If $u$ is bounded with $|u|\leqslant M$ then for each $y$ with $|y-x|^{2}>4\varepsilon M:=\delta^{2}$ one has
$\left(u(y)-\frac{1}{2\varepsilon}|y-x|^{2}\right)-u(x)<M-\frac{1}{2\varepsilon}4\varepsilon M+M=0;$
@@
$\sup_{y\in X\cap B_{\delta}(x)^{\mathsf{c}}}\left(u(y)-\frac{1}{2\varepsilon}|y-x|^{2}\right)\leqslant u(x)\leqslant u^{\varepsilon}(x).$
Hence, another expression for the supconvolution is given by
@@
Property (iv) is crucial for the quasi-convex approximation technique described in the Part II on nonlinear potential theoriesdetermined by a subequation $\mathcal{F}$. In particular, property (iv) ensures that the supconvolution of an $\mathcal{F}$-subharmonic function remains $\mathcal{F}$-subharmonic on a suitably shrunk domain (cf. Proposition 8.8).
A finer result on the correspondence between upper contact points and associated upper contact jets of a function $u$ with those of its supconvolutions $u^{\varepsilon}$ is the following, which Crandall–Ishii–Lions [13] call the magic property of the supconvolution.
@@
Let $u\in\operatorname{USC}({\mathbb{R}}^{n})$ be bounded above and, for $\varepsilon>0$, let $u^{\varepsilon}$ be its $\varepsilon$-supconvolution. If $(p,A)$ is an upper contact jet for $u^{\varepsilon}$ at $x$, then $(p,A)$ is also an upper contact jet for $u$ at $x+\varepsilon p$.
@@
which holds for every $z\in{\mathbb{R}}^{n}$ and every $y\in B_{\rho}(x)$.
By choosing $z=y-x+\xi$ in (4.15) and using the fact that $y\in B_{\rho}(x)$ is arbitrary, one has
@@
Since along a sequence $0>t_{n}\nearrow 0^{-}$ the left-hand side of the above inequality (4.16) is eventually negative when $\xi\neq x+\varepsilon p$, the only possibility for that to hold is that
@@
this implies that $u(x_{j}+\varepsilon p_{j})\to u(0)$, and the proof is concluded. ∎
@@ 4.3. The Crandall–Ishii–Lions Theorem on Sums
We will rely heavily on the properties the quasi-convex approximation via supconvolution contained in Theorem 4.9 in Part II of this paper which treats subharmonics in nonlinear potential theory. Moreover, as is well known, the quasi-convex approximation of semicontinuous functions is also a key ingredient in the theory of viscosity solutions for nonlinear PDEs. In particular, together with Jensen’s Lemma 3.6, one is able to prove the important Theorem on Sums (see the Appendix of [13]). As an application of what has been developed thus far, we give a short proof of that celebrated theorem. From our point of view, the idea is to use Proposition 4.11 to extend the quasi-convex version in Corollary 2.19 to upper semicontinuous functions.
Theorem 4.12 (On Sums; [13, Theorem 3.2]).
@@
[…]
@@
Finally, in addition to the reductions listed above, for the proof (of the theorem with these reductions) it suffices to consider a pair of functions; that is, $k=2$, as noted in [13] when they declare that the formulation of Theorem 4.12 is “formulated in a useful but distracting generality”. Taking these considerations together, one arrives at the following statement, which is also rephrased in the formalism of upper contact jets.
@@
Let $u,v\in\operatorname{USC}({\mathbb{R}}^{n})$, bounded above, with $u(0)=0=v(0)$. Define
@@
Fix $\varepsilon>0$. Easy calculations exploiting the Cauchy–Schwarz inequality yield [^7] For the sake of completeness, here are the calculations (recall that $A$ is symmetric): $\begin{split}0&\leqslant\Big{(}\sqrt{\varepsilon}|A\xi|-\frac{1}{\sqrt{\varepsilon}}|z-\xi|\Big{)}^{2}\\ &\leqslant\varepsilon|A\xi|^{2}-2\langle A\xi,z-\xi\rangle+\frac{1}{\varepsilon}|z-\xi|^{2}\\ &=\langle A\xi,\xi\rangle+\varepsilon\langle A\xi,A\xi\rangle+\frac{1}{\varepsilon}|z-\xi|^{2}+\langle A(z-\xi),z-\xi\rangle-\langle Az,z\rangle\\ &\leqslant\langle(A+\varepsilon A^{2})\xi,\xi\rangle+\Big{(}\frac{1}{\varepsilon}+\lVert A\rVert\Big{)}|z-\xi|^{2}-\langle Az,z\rangle.\end{split}$
$\langle Az,z\rangle\leqslant\langle(A+\varepsilon A^{2}),\xi\rangle+\lambda|z-\xi|^{2},\qquad\forall z,\xi\in{\mathbb{R}}^{2n},\ \lambda\vcentcolon=\frac{1}{\varepsilon}+\|A\|.$
Therefore
$w(z)-\frac{\lambda}{2}|z-\xi|^{2}\leqslant Q_{A+\varepsilon A^{2}}(\xi),$
hence, on taking the supremum over $z\in{\mathbb{R}}^{2n}$ and writing $\xi=(x,y)$, we have $(0,A+\varepsilon A^{2})$ is an upper contact jet at 00 for the supconvolution
@@
Indeed, $u^{1/\lambda}(0)\geqslant u(0)=0$ by definition, and the same holds for $v$, while $w^{1/\lambda}(0)\leqslant 0$, and thus $u^{1/\lambda}(0)=0=v^{1/\lambda}(0)$. Now, by Corollary 2.19 we know that there exist jets $(0,A_{i})$ which are the limits of upper contact jets for the $\lambda$-quasi-convex functions $u^{1/\lambda},v^{1/\lambda}$ at some points (of second-order differentiability) converging to 00 and that the $A_{i}$’s satisfy (4.19). The conclusion now follows from Proposition 4.11. ∎
@@
In view of Corollary 2.19 and the proof of Proposition 4.11, we can also say that one may choose the converging sequences of upper contact jets in Theorem 4.15 to be $(Du^{1/\lambda}(x_{j}),D^{2}u^{1/\lambda}(x_{j}))$ and $(Dv^{1/\lambda}(y_{j}),D^{2}v^{1/\lambda}(y_{j}))$, for some sequences $\{x_{j}\},\{y_{j}\}$ of points at which the supconvolutions $u^{1/\lambda},v^{1/\lambda}$ are twice differentiable.
Part II. a
\@afterheading
function $\varphi$ which is $C^{2}$ near $x\in X$ is said to be a $C^{2}$ upper test function for $u$ at $x$ if

### v3

@@ 4.2. Semiconvex approximation via sup-convolution
Let $X\subset{\mathbb{R}}^{n}$ and let $u\in\operatorname{USC}(X)$ be bounded above. For each $\varepsilon>0$ the $\varepsilon$-sup-convolution of $u$ the function defined by
@@
Before proving the results we need, here are four noteworthy facts.
@@
One can visualize how a function is transformed by the sup-convolution operation by noticing that $u^{\varepsilon}$ is the upper envelope of all quadratic functions of radius $-2\varepsilon$ with vertex at some point of the graph of $u$ [^1] By upper envelope of a family of functions $\mathscr{F}$ we mean the function $g\vcentcolon=\sup_{f\in\mathscr{F}}f$. Also note that a quadratic function with negative radius is well-defined: it just opens downwards. Also, notice that, even if we decided to “preserve” the domain of $u$ in Definition 4.5 by defining its sup-convolution only on $X$, $u^{\varepsilon}$ is in fact well-defined by (4.4) on the whole space ${\mathbb{R}}^{n}$.
@@
Moreover, it is not difficult to show that the sup-convolution is the pointwise limit of integal convolutions (mollifications). The proof of this fact is a direct application of a general large deviations result, due to Varadhan; it is given in [8] for the analogous result concerning the inf-convolution, whose definition is given in the following remark.
@@
In Theorem 4.10 below, we prove that sup-convolutions are semiconvex functions. In the obvious manner, one can also define the $\varepsilon$-inf-convolution
$u_{\varepsilon}(x)\vcentcolon=\inf_{y\in X}\Big{(}u(y)+\frac{1}{2\varepsilon}|y-x|^{2}\Big{)},\quad\text{for each}\ x\in{\mathbb{R}}^{n}.$ (4.5)
and prove that it is semiconcave. Moreover, it is not difficult to see that if $\varepsilon$ is sufficiently small, then the $\varepsilon$-sup-convolution (resp. $\varepsilon$-inf-convolution) of a semiconcave (resp. semiconvex) function remains semiconcave (resp. semiconvex). This fact, along with the characterization given in Proposition 2.5, proves a part of a well-known result of Lasry and Lions [42] on sup-inf-convolutions being $C^{1,1}$.
Remark 4.9.
The sup-convolution and inf-convolution are also used to give representation formulas for generalized solutions to Cauchy problems for Hamilton-Jacobi equations on ${\mathbb{R}}^{n}$ by way of the Hopf–Lax formula. This was first shown in [37] and extended to the viscosity setting in [3], with many subsequent developments. The simplest example is that, when $X={\mathbb{R}}^{n}$, the function $u(x,\varepsilon)\vcentcolon=u_{\varepsilon}(x)$ defined in (4.5) is the solution at time $t=\varepsilon$ of
$\partial_{t}u+\frac{1}{2}|Du|^{2}=0\quad\text{with}\quad u(x,0)=u(x).$
This helps to understand the regularizing effect of the Hamilton–Jacobi equation, since the inf-convolution of a semicontinuous initial datum is semiconcave.
Four important properties of the sup-convolution are gathered in the following theorem. We have already partially revealed the first one.
@@
Let $u\in\operatorname{USC}(X)$ be bounded above, let $u^{\varepsilon}$ be its $\varepsilon$-sup-convolution. The following hold:
@@
    $u^{\varepsilon}$ is $\frac{1}{\varepsilon}$-semiconvex on ${\mathbb{R}}^{n}$;
@@
    on $X^{\delta}\vcentcolon=\{x\in X:\ d(x,\partial X)>\delta\}$.
@@
(i)  It suffices to note that
$u^{\varepsilon}+\frac{1}{2\varepsilon}|\cdot|^{2}=\sup_{y\in{\mathbb{R}}^{n}}\bigg{(}u(y)-\frac{1}{2\varepsilon}|y|^{2}+\frac{1}{\varepsilon}\langle y,\cdot\rangle\bigg{)}$
@@
(ii)  For each $x\in X$ fixed, one has
$u^{\varepsilon}(x)=\sup_{y\in X}v_{x}(y),\quad\text{where}\quad v_{x}(y)\vcentcolon=u(y)-\frac{1}{2\varepsilon}|y-x|^{2}.$ (4.7)
Each function $v_{x}$ is upper semicontinuous on $X$ since it is the sum of upper semicontinuous functions. As $X$ is closed, extend each $v_{x}$ to be the upper semicontinuous function $\hat{v}_{x}$ defined on all of ${\mathbb{R}}^{n}$ by assigning the value $\infty$- ∞ on $X^{\mathsf{c}}$ (see Remark 4.2(a)). Notice that with this extension, one can rewrite (4.7) as
$u^{\varepsilon}(x)=\sup_{y\in{\mathbb{R}}^{n}}\hat{v}_{x}(y)\quad\forall\,x\in{\mathbb{R}}^{n}.$ (4.8)
Moreover, since $\hat{v}_{x}\equiv-\infty$ on $X^{\mathsf{c}}$ and is defined by (4.7) on $X$ (where $u$ is bounded above), one has that $\hat{v}_{x}$ is anti-coercive. Hence, for each $x\in{\mathbb{R}}^{n}$ and $\varepsilon>0$ fixed, applying Proposition 4.3(d) to $\hat{v}_{x}$ gives the existence of a point $\xi=\xi(x,\varepsilon)\in{\mathbb{R}}^{n}$ which realizes $u^{\varepsilon}(x)$ given by (4.8).
(iii)  Since $y=x\in X$ competes in the definition (4.4) of $u^{\varepsilon}(x)$ and since the penalization term in (4.4) is non positive, one has
$u(x)\leqslant u^{\varepsilon^{\prime}}(x)\leqslant u^{\varepsilon}(x)\quad\text{whenever $\varepsilon^{\prime}\leqslant\varepsilon$},\quad\forall\,x\in X.$ (4.9)
@@
where we can write “$\max$” in place of “$\sup$” since both $u$ and its quadratic penalizations are upper semicontinuous and the closed balls are compact. It is easy to see that for each fixed $x$ we have $\rho(x,\varepsilon)\searrow 0$ as $\varepsilon\searrow 0$; indeed, if not, then for all $\varepsilon$ sufficiently small, $\arg\max_{y\in X}\left(u-\frac{1}{2\varepsilon}|\cdot-\,x|^{2}\right)\subset B_{r}(x)^{\mathsf{c}}$, for some $r=r(x)>0$ independent of $\varepsilon$, yielding $u^{\varepsilon}(x)\leqslant\sup_{B_{r}(x)}u-\frac{r^{2}}{2\varepsilon}<u(x)$ if $\varepsilon$ is sufficiently small, which contradicts (4.9). Hence, the rightmost term in (4.10) converges to $u(x)$ as $\varepsilon\searrow 0$, [^2] The term converges to $\max\{u(x),\limsup_{y\to x}u(y)\}$, where $\limsup_{y\to x}u(y)\leqslant u(x)$ by the upper semicontinuity of $u$. yielding $u^{\varepsilon}(x)\searrow$ $u(x)$.
@@
(iv)  If $u$ is bounded with $|u|\leqslant M$ then for each $y$ such that $|y-x|^{2}>4\varepsilon M=\vcentcolon\delta^{2}$ one has
$u(y)-\frac{1}{2\varepsilon}|y-x|^{2}-u(x)<M-\frac{1}{2\varepsilon}4\varepsilon M+M=0;$
@@
$\sup_{y\in X\cap B_{\delta}(x)^{\mathsf{c}}}\Bigr{(}u(y)-\frac{1}{2\varepsilon}|y-x|^{2}\Bigr{)}\leqslant u(x)\leqslant u^{\varepsilon}(x).$
Hence, another expression for the sup-convolution is given by
@@
[…]
A finer result on the correspondence between upper contact points and associated upper contact jets of a function $u$ with those of its sup-convolutions $u^{\varepsilon}$ is the following, which Crandall–Ishii–Lions [14, Lemma A.5] call magical properties of the sup-convolution.
@@
Let $u\in\operatorname{USC}({\mathbb{R}}^{n})$ be bounded above and, for $\varepsilon>0$, let $u^{\varepsilon}$ be its $\varepsilon$-sup-convolution. If $(p,A)$ is an upper contact jet for $u^{\varepsilon}$ at $x$, then $(p,A)$ is also an upper contact jet for $u$ at $x+\varepsilon p$.
@@
which holds for every $z\in{\mathbb{R}}^{n}$ and every $y\in B_{\rho}(x)$. Then, by choosing $z=y-x+\xi$ in (4.13) and using the fact that $y\in B_{\rho}(x)$ is arbitrary, one has
@@
Since along a sequence $0>t_{n}\to 0^{-}$ the left-hand side of the above inequality (4.14) is eventually negative when $\xi\neq x+\varepsilon p$, the only possibility for that to hold is that
@@
this implies that $u(x_{j}+\varepsilon p_{j})\to u(0)$, which completes the proof. ∎
@@ 4.3. The Crandall–Ishii–Lions Theorem on Sums
The properties of the semiconvex approximation via sup-convolution contained in Theorem 4.10 will be used extensively in Part II, when treating subharmonics in general potential theories. Moreover, as is well known, the semiconvex approximation of semicontinuous functions is also a key ingredient in the theory of viscosity solutions for nonlinear PDEs. In particular, together with Jensen’s Lemma 3.6, one is able to prove the important Theorem on Sums (see the Appendix of [14]). As an application of what has been developed thus far, we give a short proof of that celebrated theorem. From our point of view, the idea is to use Proposition 4.12 to extend the semiconvex version of the Theorem on Sums (as discussed in Corollary 2.19) to the conventional upper semicontinuous version.
We begin by recalling the statement of the general version of the conventional result.
Theorem 4.13 (On Sums; [14, Theorem 3.2]).
@@
$w(x)\vcentcolon=u_{1}(x_{1})+\cdots+u_{k}(x_{k})\quad\text{for}\ x=(x_{1},\dots,x_{k})\in\mathscr{O},$
@@
Finally, in addition to the reductions listed above, for the proof (of the theorem with these reductions) it suffices to consider a pair of functions; that is, $k=2$, as noted in [14] when they declare that Theorem 4.13 is “formulated in a useful but distracting generality”. Taking these considerations together, one arrives at the following statement, which we also rephrased in the formalism of upper contact jets.
@@
Let $u,v\in\operatorname{USC}({\mathbb{R}}^{n})$ be bounded above and with $u(0)=0=v(0)$. Define
@@
Fix $\varepsilon>0$. Easy calculations exploiting the Cauchy–Schwarz inequality yield
$\begin{split}0&\leqslant\Big{(}\sqrt{\varepsilon}|A\xi|-\frac{1}{\sqrt{\varepsilon}}|z-\xi|\Big{)}^{2}\\ &\leqslant\varepsilon|A\xi|^{2}-2\langle A\xi,z-\xi\rangle+\frac{1}{\varepsilon}|z-\xi|^{2}\\ &=\langle A\xi,\xi\rangle+\varepsilon\langle A\xi,A\xi\rangle+\frac{1}{\varepsilon}|z-\xi|^{2}+\langle A(z-\xi),z-\xi\rangle-\langle Az,z\rangle\\ &\leqslant\langle(A+\varepsilon A^{2})\xi,\xi\rangle+\Big{(}\frac{1}{\varepsilon}+\lVert A\rVert\Big{)}|z-\xi|^{2}-\langle Az,z\rangle;\end{split}$
that is,
$\langle Az,z\rangle\leqslant\langle(A+\varepsilon A^{2}),\xi\rangle+\lambda|z-\xi|^{2}\qquad\forall\,z,\xi\in{\mathbb{R}}^{2n},\ \ \text{with}\ \lambda\vcentcolon=\frac{1}{\varepsilon}+\|A\|.$
Therefore, we have
$w(z)-\frac{\lambda}{2}|z-\xi|^{2}\leqslant Q_{A+\varepsilon A^{2}}(\xi).$
Now, by taking the supremum over $z\in{\mathbb{R}}^{2n}$ and by writing $\xi=(x,y)$, we have that $(0,A+\varepsilon A^{2})$ is an upper contact jet at 00 for the sup-convolution
@@
Indeed, $u^{1/\lambda}(0)\geqslant u(0)=0$ by definition, and the same holds for $v$, while $w^{1/\lambda}(0)\leqslant 0$, and thus $u^{1/\lambda}(0)=0=v^{1/\lambda}(0)$. Now, by Corollary 2.19 we know that there exist jets $(0,A_{i})$ which are the limits of upper contact jets for the $\lambda$-semiconvex functions $u^{1/\lambda},v^{1/\lambda}$ at some points (of second-order differentiability) converging to 00 and that the $A_{i}$’s satisfy (4.17). The conclusion now follows from Proposition 4.12. ∎
@@
In view of Corollary 2.19 and the proof of Proposition 4.12, we can also say that one may choose the converging sequences of upper contact jets in Theorem 4.16 to be $(Du^{1/\lambda}(x_{j}),D^{2}u^{1/\lambda}(x_{j}))$ and $(Dv^{1/\lambda}(y_{j}),D^{2}v^{1/\lambda}(y_{j}))$, for some sequences $\{x_{j}\}_{j\in\mathbb{N}},\{y_{j}\}_{j\in\mathbb{N}}$ of points at which the sup-convolutions $u^{1/\lambda},v^{1/\lambda}$ are twice differentiable.
Part II General potential–theoretic analysis
This second part consists of four chapters and is dedicated to general nonlinear second order potential theories and the fundamental role that semiconvex functions play in such theories. The presentation will involve four stages which include a derivation of the minimal conditions that a subequation constraint set should satisfy in order to have a robust potential theory of subharmonic functions. The fundamental roles of duality and monotonicity will be discussed, along with the main tools in the viscosity theory of subharmonics and the essential role that semiconvex functions play in that theory. The validity of the comparison principle will be examined in two distinct regimes of sufficient monotonicity and minimal monotonicity. The final brief chapter shows that the notion of semiconvexity is preserved by smooth changes of coordinates, which is the key point in order to transport semiconvex analysis to manifolds. It is included here because we make use of potential theoretic arguments.
Chapter 5 General potential theories
In this chapter we will discuss the main objects in general potential theories; that is, subequation constraint sets and their subharmonics. Both classical and viscosity formulations will be discussed as well as the relations between the two notions. A few representative potential theories will be presented; namely classical subharmonics, convex and semiconvex functions.
5.1. Subequation constraint sets and their subharmonics
[…]
Potential theories associated to subequations are interesting from two different points of view. First, there are many interesting geometric contexts with naturally associated potential theories as reviewed in [35, Section 1.1]. Second, subequation constraint sets and their associated potential theory can arise naturally in the study of a given (weakly elliptic) PDE, which we recall now for the benefit of those familiar with the standard viscosity theory of PDEs but not with the corresponding general potential theoretic context.
Suppose we are given an operator $F\in C(X\times{\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n))$ on some open set $X\subset{\mathbb{R}}^{n}$ which determines a PDE
$F(x,u(x),Du(x),D^{2}u(x))=0,\quad x\in X.$ (5.1)
A classical subsolution of (5.1) is a function $u\in C^{2}(X)$ which satisfies
$F(x,u(x),Du(x),D^{2}u(x))\geqslant 0\quad\forall x\in X.$ (5.2)
[…]
:=$\mathcal{F}\vcentcolon=\{(x,p,r,A)\in\mathcal{J}^{2}(X):F(x,r,p,A)\geqslant 0\},$ (5.3)
[…]
$J^{2}_{x}u\vcentcolon=(u(x),Du(x),D^{2}u(x))\in\mathcal{F}_{x}\qquad\forall x\in X,$
where
$\mathcal{F}_{x}\vcentcolon=\left\{(r,p,A)\in{\mathbb{R}}\times{\mathbb{R}}^{n}\times\mathcal{S}(n):\ (x,r,p,A)\in\mathcal{F}\right\}$
is the fiber of $\mathcal{F}$ over $x$. One extends the notions of subsolutions and subharmonics to upper semicontinuous functions by using a pointwise viscosity formulation in terms of $2$-jets of spaces of upper contact test functions; that is, spaces of upper contact jets. Under minimal conditions on $\mathcal{F}$, there are many equivalent formulations as will be discussed below in Remark 5.4.
Definition 5.1.
Given a function $u\in\operatorname{USC}(X)$:
  • (a)
    a function $\varphi$ which is $C^{2}$ near $x\in X$ is said to be a $C^{2}$ upper test function for $u$ at $x$ if